Please show all steps. Thanks

Please show all steps. Thanks
Answer #1

4) tem is system non homogeneous У? 2 81 - 92 у? 91 + чу. - 6x — 1ъ Зя - 2. Differentiating o urrit. x we have - 9 - 29. -4 y, - 6y,! +9y 18% +.15 Auxiliary equation for the homogeneous part. is m²-6m + 9 (m-3)2 =0 m = 3,3 is 9 e 34 que se be so— Therefore 1 ts) - Сте“ +4, sre* + 2 +-2 ц, (м) - 2.сезм + 2C, xe3+ + c = 3х -2. and Зc, e3 эг (e3х – 30xe3x 2. З, (x) - (-[ 3 (7)(s A E 4 Hace A = 6 and IAI = 8+1=9 eigenealue of a werthe multiplicity2 3 18 an Roots of characteriatic equation aPLEASE GIVE A RATING IF THIS WAS HELPFUL FOR YOU.